OpenStudy (anonymous):

"Gina rented a moon walk for 5 hours at a total cost of $235. Lori rented the same moon walk for 3 hours for a total cost of $165. The cost of renting is based on a flat set-up fee + a rental rate per hour. How much is the set up fee? How much is the hourly rental fee?" Please explain. I've tried everything I can think of!

OpenStudy (anonymous):

you can look at the information and see there will be two functions: we can call them Gina Lori

OpenStudy (anonymous):

each function will have a flat set up fee plus a rental rate

OpenStudy (anonymous):

you know?

OpenStudy (anonymous):

yes, but how do you solve?

OpenStudy (anonymous):

x+5y=@235 x+3y=165

OpenStudy (anonymous):

solve the bottom equation for x x=165-3y substitute into the first equation to solve for y 165-3y+5y=235

OpenStudy (anonymous):

what do you get for y

OpenStudy (anonymous):

@Kiralee

OpenStudy (anonymous):

y=35

OpenStudy (anonymous):

wth? did that and it didn't work.... thank you

OpenStudy (anonymous):

it didn't work? @Kiralee

OpenStudy (anonymous):

no, it did this time, but i did the same thing by myself 3 times before asking, got the same answers, but when i went to check it it was somehow off every time

OpenStudy (anonymous):

its always easier to start with the lesser of the two equations

OpenStudy (anonymous):

substituted into the larger of the two